# Where Are We? Evaluating LLM Performance on African Languages

### v1

@@ Abstract
Africa’s rich linguistic heritage remains underrepresented in NLP, largely due to historical policies that favor foreign languages and create significant data inequities. In this paper, we integrate theoretical insights on Africa’s language landscape with an empirical evaluation using Sahara— a comprehensive benchmark curated from large-scale, publicly accessible datasets capturing the continent’s linguistic diversity. By systematically assessing the performance of leading large language models (LLMs) on Sahara, we demonstrate how policy-induced data variations directly impact model effectiveness across African languages. Our findings reveal that while a few languages perform reasonably well, many Indigenous languages remain marginalized due to sparse data. Leveraging these insights, we offer actionable recommendations for policy reforms and inclusive data practices. Overall, our work underscores the urgent need for a dual approach—combining theoretical understanding with empirical evaluation—to foster linguistic diversity in AI for African communities.
@@ 1 Introduction
The digital landscape of Africa is as diverse as its cultures and languages, yet the continent’s rich linguistic heritage remains largely underrepresented in NLP research. Despite significant transformations over the past decade driven by advances in large language models (LLMs) and a growing interest in linguistic diversity, most efforts have concentrated on a handful of widely spoken languages such as Swahili, Afrikaans, and Hausa, while many Indigenous languages continue to be sidelined due to the paucity of accessible, high-quality datasets. This oversight, deeply rooted in historical language policies that have neglected robust data collection and accessibility, has far-reaching implications for the development of language technology across Africa Eberhard et al. (2021). Recent initiatives, including the work of Masakhane Dione et al. (2023b); Adelani et al. (2021c, 2023, 2024b); Bandarkar et al. (2024a) and the efforts of platforms like HuggingFace, have begun to address these challenges by curating language-specific datasets and adapting transfer learning techniques for low-resource African languages. Models such as Cheetah Adebara et al. (2024a), Serengeti Adebara et al. (2023), Toucan Elmadany et al. (2024), Afri-XLMR Alabi et al. (2022), AfriTeVa Oladipo et al. (2023); Jude Ogundepo et al. (2022), mBERT Devlin et al. (2018), XLM-R Conneau et al. (2020), and LLaMA have leveraged multilingual data to enhance performance on these languages, yet progress remains uneven—highlighting that the advancements achieved are primarily concentrated on a select few.
In this study, we undertake a comprehensive empirical evaluation of leading large language models (LLMs) on an extensive benchmark that we collect using mostly existing datasets to unravel how current African language policies, by dictating data availability and accessibility, directly shape model performance across the continent’s diverse linguistic spectrum. By systematically analyzing the distribution of existing datasets across languages and performance patterns among various LLMs, we not only pinpoint which languages benefit from ample data resources but also illuminate the underlying factors that contribute to their relative success. Our investigation reveals that the disparities in NLP outcomes are closely tied to policy-driven data inequities, offering concrete evidence of the need for more inclusive, forward-thinking language policies. These insights provide actionable recommendations for enhancing dataset creation and model training, ultimately aiming to bridge the digital divide and foster linguistic diversity in artificial intelligence for African communities.
@@ 2 Language Policies in Africa
(a) Language policies for education across Africa.
(b) Official national language policies across Africa.
(c) Official regional language policies across Africa. None means no additional policy is available on a regional level.
(d) We collect  Saharato empirically evaluate LLM performance on African languages, allowing us to demonstrate with evidence how current language policies directly impact progress in the field.
Figure 1: Maps of Africa showing the languages covered in this work and the language policies across Africa. The term Both in maps (a), (b), and (c) refer to Indigenous and foreign languages combined.
Across Africa, the predominant response to the continent’s multilingual landscape has been the adoption of a foreign language for official functions. In some instances, select Indigenous African languages receive official recognition at regional or national levels or within educational contexts Petzell (2012); Foster (2021); Ouane and Glanz (2010). Figures 1(a), 1(b), and 1(c) illustrate the distribution of language policies across the continent. [^1] We designed the maps in these figures based on information derived from Ethnologue. We checked the national, regional, institutional, and educational languages for each country. For example, in Nigeria, English is the official language, with only three out of 512 Indigenous languages recognized at the regional level. Similarly, Ghana designates English as its official language while also recognizing ten of its 73 Indigenous languages for institutional use. In Tanzania, Swahili is the sole Indigenous official language among 118 languages, alongside English. Kenya grants official status to $12$ of its $61$ languages, while South Africa recognizes $12$ out of $20$ Indigenous languages as institutional languages Adebara and Abdul-Mageed (2022).
@@ 3 Related Work
The development of NLP for African languages is hindered by data scarcity, policy-induced disparities, and suboptimal model performance. Recent initiatives have sought to mitigate these challenges by curating diverse datasets, rigorously evaluating multilingual models, and advocating for more inclusive language policies. For example, several initiatives have contributed to expanding datasets for African Yimam et al. (2020); Muhammad et al. (2022); Aliyu et al. (2022); Muhammad et al. (2023b, a); Ilevbare et al. (2024); Elmadany et al. (2024). Some notable ones like Adelani et al. (2021c, 2022b) introduced MasakhaNER, a named entity recognition dataset for 10 African languages, highlighting the limitations in existing resources. Similarly, IROKOBench Adelani et al. (2024b) evaluated large language models (LLMs) on African languages, revealing stark performance gaps between high- and low-resource languages. Adebara et al. (2022) developed AfroLID, a neural language identification model covering 517 African languages, showing that most languages lack NLP datasets beyond language identification.
@@
Our objective is to create a comprehensive benchmark for African NLP that enables (i) analysis of existing language resources and (ii) the assessment of language models and tracking of the progress of African NLP. To achieve this goal, we develop Sahara, adhering to several key design principles that we will now elucidate. Sahara establishes a comprehensive and adaptable benchmark for African NLP using publicly available datasets. Figure 1(d) shows coverage of Sahara.
@@
Cluster Identifier Metric SLMs LLMs Closed LLMs
S1 S2 S3 S4 L1 L2 L3 L4 L5 L6 L7 CL1 CL2
Classification xnli Acc. 32.04 44.44 40.56 47.78 54.26 46.48 61.67 68.52 65.00 47.04 42.96 74.41 82.65
langid Acc. 0.24 0.59 0.70 1.89 1.11 2.34 2.21 2.45 8.69 17.64 15.45 20.56 13.34
news Acc. 13.76 12.96 2.00 2.77 42.86 2.66 18.52 54.16 27.27 75.83 7.85 58.33 67.85
sentiment Acc. 45.65 22.25 30.30 14.58 45.00 7.14 33.92 25.33 49.25 69.69 25.33 56.33 56.45
topic Acc. 36.10 34.28 17.27 15.80 57.77 10.70 38.09 42.63 78.43 67.43 15.38 57.70 79.99
Avg. 25.56 22.90 18.17 15.36 40.20 13.86 30.88 38.62 45.73 55.53 21.39 53.47 60.06
Generation mt Bleu 1.82 7.31 0.04 3.66 6.18 5.26 6.91 5.03 17.48 13.80 5.57 9.99 18.34
paraphrase Bleu 13.17 16.27 10.05 7.12 19.14 8.05 6.14 9.83 24.67 37.37 8.22 17.25 37.00
summary Rouge 7.55 9.98 2.00 9.13 11.18 9.56 2.00 14.69 13.24 11.01 14.45 9.10 12.35
title-gen Bleu 0.32 1.95 0.84 2.09 4.01 0.26 1.68 3.95 14.68 4.81 2.86 4.28 13.55
Avg. 5.72 8.88 3.23 5.50 10.13 5.78 4.18 8.38 17.52 16.75 7.78 10.16 20.31
MCCR mmlu Acc. 27.59 27.59 36.30 38.70 61.85 34.26 69.26 76.67 71.11 61.85 41.48 70.00 75.00
mgsm ExctM 1.94 1.94 1.48 5.19 8.24 1.11 12.04 15.56 17.04 10.00 5.19 16.98 22.34
rdng-comp Acc. 24.65 33.47 30.83 30.90 40.28 33.46 47.01 49.23 51.03 40.40 38.33 78.54 85.19
squad-qa F1 59.36 69.46 42.80 46.90 73.06 29.68 43.67 79.43 87.47 77.35 47.58 58.67 55.99
Avg. 28.39 33.12 27.85 30.42 45.86 24.63 43.00 55.22 56.66 45.90 33.15 56.05 59.63
Tokens ner Acc. 17.77 27.90 19.65 16.06 29.37 0.56 22.23 49.71 37.91 42.82 7.20 46.70 16.94
chunking Acc. 5.02 11.02 6.92 4.90 21.61 0.05 12.57 18.15 15.15 19.42 2.82 26.60 18.21
pos Acc. 6.34 6.96 2.80 5.12 11.13 1.31 9.05 15.09 0.06 10.10 3.46 11.02 11.04
Avg. 9.71 15.29 9.79 8.69 20.70 0.64 4.62 27.65 17.71 24.11 4.49 28.11 15.40
Overall Avg. 17.34 20.05 14.76 14.99 29.22 11.23 23.17 32.47 34.40 35.57 16.70 36.94 38.85
Table 2: Few-shot evaluation across task datasets. The highest score for each individual task is in bold green, while the best average score across all tasks is in bold orange. S1Llama3.2 (1B) S2 Llama3.2 (3B) S3 Gemma2 (2B) S4 Phi3.5 (3.8B) L1LLama3.1 (8B) L2 Aya23 (8B) L3 Gemma2 (9B) L4 Gemma2 (27B) L5 Llama3.1 (70B) L6 CommadR+ (104B) L7 Aya23 (35B) CL1 GPT4o CL1 Claude-3.5
@@ 5.1 Experimental Setup
To provide an overview of the state of African NLP, we evaluated the performance of popular instruction-based LLMs in a few-shot setting using the Sahara benchmark. The models tested include Phi-3.5 Abdin et al. (2024), Gemma-2 Team et al. (2024), Llama-3.1 Dubey et al. (2024), Llama-3.2, Aya-23 Aryabumi et al. (2024), and command-r-plus from open LLMs, as well as GPT-4o and Claude-3.5-Sonnet from closed LLMs. We also considered variations of these models based on their size, categorizing the LM as small, SLM for those with fewer than 8B parameters and large, LLM for those with more than 8B parameters to analyze the effect of model size on performance. For clearer visualization, we aggregated the test datasets for each task cluster identified in Appendix A. We randomly chose a sample of $600$ data points for each task for a few-shot evaluation. The evaluation used well-crafted prompts that are both concise and straightforward, with examples of these prompts provided in Appendix C1. For tasks available in the lm-eval tool Gao et al. (2024), we use the tool for evaluating the open-source LMs.

### v2

# Evaluating LLM Performance on African Languages

@@ Abstract
Africa’s rich linguistic heritage remains underrepresented in NLP, largely due to historical policies that favor foreign languages and create significant data inequities. In this paper, we integrate theoretical insights on Africa’s language landscape with an empirical evaluation using Sahara— a comprehensive benchmark we curate from large-scale, publicly accessible datasets capturing the continent’s linguistic diversity. By systematically assessing the performance of leading large language models (LLMs) on Sahara, we demonstrate how policy-induced data variations directly impact model effectiveness across African languages. [^1] Sahara is publicly available at https://github.com/UBC-NLP/sahara. Our findings reveal that while models perform reasonably well on few languages, many Indigenous languages remain marginalized due to sparse data. Leveraging these insights, we offer actionable recommendations for policy reforms and inclusive data practices. Overall, our work underscores the urgent need for a dual approach—combining theoretical understanding with empirical evaluation—to foster linguistic diversity in AI for African communities.
{strip}
[Uncaptioned image] (a) Language policies for education across Africa. [Uncaptioned image] (b) Official national language policies across Africa.
[Uncaptioned image] (c) Official regional language policies across Africa. None means no additional policy is available on a regional level. [Uncaptioned image] (d) We collect Sahara to empirically evaluate LLM performance on African languages, allowing us to demonstrate with evidence how current language policies directly impact progress in the field.
Figure 1: Maps of Africa showing the languages covered in this work and the language policies across the continent. The term Indigenous language is broadly defined here as one that is ‘native’ to the area Walsh (2005). The term Both in maps (a), (b), and (c) refers to Indigenous and foreign languages combined. Knowledge of what is an Indigenous language is based exclusively on Ethnologue (https://www.ethnologue.com).
@@ 1 Introduction
The digital landscape of Africa is as diverse as its cultures and languages, yet the continent’s rich linguistic heritage remains largely underrepresented in NLP research. Despite significant transformations over the past decade driven by advances in large language models (LLMs) and a growing interest in linguistic diversity, most efforts have concentrated on a handful of widely spoken languages such as Swahili, Afrikaans, and Hausa, while many Indigenous languages continue to be sidelined due to the paucity of accessible, high-quality datasets. This oversight, deeply rooted in historical language policies that have neglected robust data collection and accessibility, has far-reaching implications for the development of language technology across Africa Eberhard et al. (2021). Recent initiatives, including the work of Masakhane Dione et al. (2023b); Adelani et al. (2021c, 2023, 2024b); Bandarkar et al. (2024a) and the efforts of platforms like HuggingFace, have begun to address these challenges by curating language-specific datasets and adapting transfer learning techniques for low-resource African languages. Models such as Cheetah Adebara et al. (2024a), Serengeti Adebara et al. (2023), Toucan Elmadany et al. (2024), Afri-XLMR Alabi et al. (2022), AfriTeVa Oladipo et al. (2023); Jude Ogundepo et al. (2022), mBERT Devlin et al. (2018), XLM-R Conneau et al. (2020), and LLaMA Touvron et al. (2023) have leveraged multilingual data to enhance performance on these languages, yet progress remains uneven—highlighting that the advancements achieved are primarily concentrated on a select few.
In this study, we undertake a comprehensive empirical evaluation of leading LLMs on an extensive benchmark that we collect using mostly existing datasets to unravel how current African language policies, by dictating data availability and accessibility, directly shape model performance across the continent’s diverse linguistic spectrum. By systematically analyzing the distribution of existing datasets across languages and performance patterns among various LLMs, we not only pinpoint which languages benefit from ample data resources but also illuminate the underlying factors that contribute to their relative success. Our investigation reveals that the disparities in NLP outcomes are closely tied to policy-driven data inequities, offering concrete evidence of the need for more inclusive, forward-thinking language policies. These insights provide actionable recommendations for enhancing dataset creation and model training, ultimately aiming to bridge the digital divide and foster linguistic diversity in artificial intelligence for African communities.
@@ 2 Language Policies in Africa
Across Africa, the predominant response to the continent’s multilingual landscape has been the adoption of a foreign language for official functions. However, in some instances, select Indigenous African languages receive official recognition at regional or national levels or within educational contexts Petzell (2012); Foster (2021); Ouane and Glanz (2010).  [^2] An Indigenous language is defined here as one that is native to a particular region, rather than introduced from elsewhere, in contrast to a foreign language, which may nonetheless serve as the mother tongue for certain local populations Walsh (2005). A national language refers to a language spoken broadly across an entire country. A regional language is spoken by a significant number of people within a specific area but does not have widespread national usage. An educational language is officially designated as a medium of instruction within the educational system. For instance in Nigeria, English is the official language, with only three out of $512$ Indigenous languages recognized at the regional level. Similarly, Ghana designates English as its official language while also recognizing ten of its $73$ Indigenous languages for institutional use. In Tanzania, Swahili is the sole Indigenous official language among $118$ languages, alongside English. Kenya grants official status to $12$ of its $61$ languages, while South Africa recognizes $12$ out of $20$ Indigenous languages as institutional languages Adebara and Abdul-Mageed (2022). Figures 1(a),  1(b),  1(c), and  1(d) illustrate the distribution of language policies across the continent.
@@ 3 Related Work
The development of NLP for African languages is hindered by data scarcity, policy-induced disparities, and suboptimal model performance. Recent initiatives have sought to mitigate these challenges by curating diverse datasets, rigorously evaluating multilingual models, and advocating for more inclusive language policies. For example, several initiatives have contributed to expanding datasets for African Yimam et al. (2020); Muhammad et al. (2022); Aliyu et al. (2022); Muhammad et al. (2023b, a); Ilevbare et al. (2024); Elmadany et al. (2024). Some notable ones like Adelani et al. (2021c, 2022b) introduced MasakhaNER, a named entity recognition dataset for 10 African languages, highlighting the limitations in existing resources. Similarly, IROKOBench Adelani et al. (2024b) evaluated LLMs on African languages, revealing stark performance gaps between high- and low-resource languages. Adebara et al. (2022) developed AfroLID, a neural language identification model covering 517 African languages, showing that most languages lack NLP datasets beyond language identification.
@@
Our objective is to create a comprehensive benchmark for African NLP that enables (i) analysis of existing language resources and (ii) the assessment of language models and tracking of the progress of African NLP. To achieve this goal, we develop Sahara, adhering to several key design principles that we will now elucidate. Sahara establishes a comprehensive and adaptable benchmark for African NLP using publicly available datasets. Figure [Uncaptioned image] Where Are We? Evaluating LLM Performance on African Languages shows coverage of Sahara.
@@
SLM LLM Closed LLM
Cluster Identifier Metric Shots S1 S2 S3 S4 S5 S6 L1 L2 L3 L4 L5 L6 L7 L8 L9 L10 L11 L12 L13 L14 L15 L16 CL1 CL2
Classification xlni Acc. 5 33.6 33.7 36.7 41.18 40.4 46.68 48.08 55.06 34.3 40.69 37.09 61.09 57.86 65.58 44.2 38.19 51.89 48.38 52.09 35.87 46.77 59.67 69 69.6
lid $F_{1}$ 10 0 0 0.87 0.26 0.85 0.77 2.82 3.6 0 1.49 2.27 1.95 4.61 4 1.81 3.34 3.26 3.07 3.3 3.05 3.32 4.57 3.73 4.48
news $F_{1}$ 3 0.86 0.35 0.15 17.54 0 6.64 21.94 0 2.06 24.06 0.12 25.83 7.83 35.56 28.03 0.61 35.26 33.31 34.81 3.45 30.46 28.66 40.09 32.19
sentiment $F_{1}$ 5 38.51 24.75 30.45 11.05 28.77 1.99 36.09 42.47 38.4 28.36 36.07 34.56 20.55 46.35 35.86 25.94 24.51 45.82 35.89 19.98 37.07 47.14 55.67 57.52
topic $F_{1}$ 3 8.73 31.62 20.3 19.79 12.46 17.15 48.11 39.35 34.95 28.55 42.12 44.3 31.8 70.73 36.47 43.74 64.89 56.48 48.73 52.09 50.96 53.16 67.93 76.57
Avg. 16.34 18.08 17.69 17.96 16.50 14.65 31.41 28.10 21.94 24.63 23.53 33.55 24.53 44.44 29.27 22.36 35.96 37.41 34.96 22.89 33.72 38.64 47.28 48.07
Generation mt_eng2xx spBleu1K 5 1.79 3.63 3.45 3.25 2.12 3.4 7.07 7.61 1.72 4.37 2.79 6.24 6.91 4.56 3.73 3.29 8.84 8.7 9.36 1.4 6.06 8.22 12.18 12.36
mt_fra2xx spBleu1K 5 0.79 0.36 0.74 1.05 1.25 1.32 0.73 0.95 0.7 1.19 0.66 2.03 1.67 2.05 1.15 0.82 1.93 1.87 1.77 1.05 1.04 2.3 3.41 3.42
mt_xx2xx spBleu1K 5 0.4 0.45 0.42 0.42 0.36 0.18 0.43 0.63 0.72 0.55 0.46 0.71 0.9 1.39 0.55 0.17 1.28 1.05 1.37 0.67 0.38 0.85 5.15 4.44
paraphrase spBleu1K 5 23.2 28.88 31.15 19.12 25.93 17.85 28.73 25.83 31.68 31.33 23.66 17.26 21.34 15.37 32.13 30.68 26.86 19.64 27.79 19.23 36.06 26.06 23.2 20.35
summary rougeL 2 10.03 3.24 4.37 10.07 0.91 13.34 15.1 10.43 1.45 12.74 5.92 16.16 15.73 17.13 13.27 4.19 16.92 17.19 16.4 15.46 14.88 16.09 5.99 16.3
title spBleu1K 2 2.14 2.81 0.25 2.3 0.03 5.78 7.48 5.78 3 5.94 1.58 7.75 7.51 8.65 6.83 3.2 11.07 10.34 8.78 7.32 6.69 8.65 13.62 9.48
Avg. 6.39 6.56 6.73 6.04 5.10 6.98 9.92 8.54 6.55 9.35 5.85 8.36 9.01 8.19 9.61 7.06 11.15 9.80 10.91 7.52 10.85 10.36 10.59 11.06
MCCR mmlu Acc. 5 24.09 28.5 28.68 27.29 28.9 30.99 31.88 40.9 26.4 44.51 22.1 42.09 44.9 60.71 51.5 24 56.51 59.3 50.42 59.81 29.8 61.91 81.6 58.6
mgsm ExactM 5 2.6 3.21 0 1.2 3.39 0.7 2.9 2.8 3.3 2.7 1.5 2.3 5.3 0.6 4.4 3 10.1 7.51 7.6 4.5 4.39 11.1 45.9 29.5
belebele Acc. 5 27.71 14.71 25.2 29.31 31.31 24.69 22.7 26.93 25.09 28.24 26.33 25.63 31.04 32.34 33.64 26.61 30.4 35.92 33.72 30.83 20.99 29.12 36.6 39.7
squad_qa $F_{1}$ 5 57.45 70.89 67.32 43.13 71.31 74.97 78.04 80.91 52.66 59.48 63.41 73.78 82.49 79.16 69.25 64.64 77.65 76.35 66.25 53.92 74.64 80.05 78.02 76.11
Avg. 27.96 29.33 30.30 25.23 33.73 32.84 33.88 37.89 26.86 33.73 28.34 35.95 40.93 43.20 39.70 29.56 43.67 44.77 39.50 37.27 32.46 45.55 60.53 50.98
Tokens ner $F_{1}$ 5 0 0.78 1.76 2.99 0.13 3.65 2.26 7.49 2.35 3.96 0.2 6.03 13.14 11.4 8.51 1.52 6.86 10.54 8.64 6.04 4.3 16.88 35.06 27.04
phrase $F_{1}$ 5 14.79 28.5 29.44 11.52 24.91 27.73 15.95 29.35 29.42 16.55 30.87 13.59 30.01 23.97 26.34 23.09 25.55 23.13 16.78 25.13 29.44 31.05 36.69 36.45
pos $F_{1}$ 5 8.83 6.73 10.56 8.04 10.3 10.46 9.13 13.2 9.81 11.68 12.52 13.86 19 13.99 18.36 11.4 14.13 15.04 11.52 20.76 13.71 27.54 62.84 38.67
Avg. 7.87 12.00 13.92 7.52 11.78 13.95 9.11 16.68 13.86 10.73 14.53 11.16 20.72 16.45 17.74 12.00 15.51 16.24 12.31 17.31 15.82 25.16 44.86 34.05
Overall 14.64 16.49 17.16 14.19 16.78 17.10 21.08 22.80 17.30 19.61 18.06 22.25 23.80 28.07 24.08 17.75 26.57 27.05 24.42 21.25 23.21 29.93 40.82 36.04
Table 2: Few-shot evaluation across different task clusters/tasks. The highest score for each individual task is in bold green, while the best average score across each task cluster (and overall score, calculated as average of task cluster averages) is in bold orange. Underline: refers to the best score across the open source/weights LLMs. Shots: number of shots. SLMs: S1 Llama3.2 (1B), S2 Llama3.2 (3B), S3 Gemma2 (2B), S4 Phi-3.5 (3.8B), S5 Phi-4 (3.8B), and S6 Gemma3 (4B). LLMs: L1 Llama3.1 (8B), L2 Gemma2 (9B), L3 Aya (8B), L4 Babel (9B), L5 Command-R-Plus-R7B (8B), L6 Gemma3 (12B), L7 Gemma2 (27B), L8 Gemma3 (27B), L9 DeepSeek-R1-Distill-Qwen (32B), L10 Aya (35B), L11 Llama3.1 (70B), L12 Llama3.3 (70B), L13 DeepSeek-R1-Distill-Qwen (70B), L14 Babel (83B), L15 Command-R-Plus (104B), L16 Command-A ( [^1] B). CLosed LLMs: CL1 Claude-4-Sonnet, CL2 GPT-4.1. Table E4 is a full-page version of this table.
@@ 5.1 Experimental Setup
To comprehensively assess the progress of African NLP, we systematically evaluate a range of widely used instruction-based LLMs in a few-shot setting on the Sahara benchmark. We include both open and closed LLMs, such as Phi-3.5 and Phi-4 Abdin et al. (2024), Gemma family Team et al. (2024) models including Gemma-2 (2, 9, 27B) and Gemma-3 (4, 12, 27B), Llama family models Dubey et al. (2024) including Llama-3.1 (8, 70B), Llama-3.2 (1, 3B), and Llama-3.3 (70B), Aya models (8, 23B) Aryabumi et al. (2024), and Command family models Cohere et al. (2025) including Command-R-Plus-R7B, Command-R-Plus (104B) and Command-A ( [^1] B), as well as proprietary models such as GPT-4.1333GPT-41.1 version: gpt-4.1-2025-04-14.  Achiam et al. (2023) and Claude-4-Sonnet [^4] Claude-4-Sonnet version: claude-4-sonnet-20250514  Anthropic (2024). We further analyze performance by categorizing models according to size: small language models (SLM, $<$8B parameters) and large language models (LLM, $\geq$8B parameters), allowing investigation of the relationship between scale and effectiveness. For robust comparison, we aggregate evaluation datasets according to task clusters described in Appendix A. From each task, we randomly sample $1,000$ data points for few-shot testing, employing well-curated, concise prompts—examples of which appear in Appendix C1.